OpenStudy (anonymous):

Which expression is not equivalent to the other expressions? -6(2x- 4) -(12x- 6) + 18 -3(4x- 3) + 15 -4(3x+ 6)

OpenStudy (jh3power):

?

OpenStudy (anonymous):

Which expression is not equivalent to the other expressions? -6(2x- 4) -(12x- 6) + 18 -3(4x- 3) + 15 -4(3x+ 6)

OpenStudy (jh3power):

-6(2x-4)=-12x+24 -(12x-6)+8=-12x+14 -3(4x-3)+15=-12x+24 -4(3x+6)=-12x-24 so..second one